HUBBLE SPACE TELESCOPE DAILY REPORT #865 PERIOD COVERED: 0800 EST Daily Status Report as of 104/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 3791 (Quasar Absorption Line Survey: Cycle 2 Observations) The Faint Object Spectrograph (Red Side) was used to make six spectral observations of the quasi-stellar object PG1538+477. A binary search was followed by a peakup. The target was found in the very last group of the peakup. The subsequent observations were made with the 0.25 X 2.0 arcsecond slit. All six observations were received and displayed in the OSS. They were reported to look okay. 1.2 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.3 Completed FGS 4725 (Scale Cycle 2 Calibration) The Fine Guidance Sensors were used to make astrometric observations to calibrate the scale in the central region of the astrometer FGS. The calibration is required to support the reduction of double star observations. Since FGS #2 is no longer available, the observations were done on the given target (Star HD155103) with one FGS and gyro guiding. The observations were executed as scheduled, and no problems were reported. 1.4 Completed HSP/PMT/VIS 4257 (Dynamics of Planetary Upper Atmosphere - Uranus, Part 2 (1083)) Observations were made using the High Speed Photometer of the Planet Uranus during an occultation by a faint star in Capricornus. These observations will help in our understanding of the planet's atmospheric radiative and dynamical processes. This event occurred close to the last quarter moon, and special arrangements had to be made to modify the lunar limit tests to allow these observations. The OSS is currently reviewing the observations, and reported initially all the observations looked okay. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 10 Successful : 10 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 0 Successful : 0 2.3 Significant HSTARs: None 2.4 Operations Note: GSFC Simulation #5 (Deployment) is currently (104/1200Z through 2400Z) being conducted in the MOR. No problems have been reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 4221 (Imaging and Spectrophotometry of Nuclear Activity in Liners (FOS 15)) 3.2 Complete PC 4140 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor (Part 2)) 3.3 Complete FGS 4031 (High Speed Astrometry - A Search for Planetary Companions to Low-Mass) 3.4 Complete FGS 3688 (Diagnostic of Early Stellar Evolution: Search for Pre-main Sequence Stars)) 3.5 Complete HSP/VIS/PMT 4257 (Dynamics of Planetary Upper Atmospheres - Part 2 (1083)) 3.6 Complete FOS/RD 4528 (SINS: The Supernova Intensive Study - SN1993J) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time.